# On the Brun–Titchmarsh theorem

### v1

@@ Acknowledgements
This work is supported in part by NSFC (No. 12025106).
2. Linear sieves
$A_{d}=\sum_{n\equiv 0\!\,(\operatorname{mod}{d})}a_{n}.$
In applications to sieve methods, it is expected that $\mathcal{A}$ equidistributes in the special arithmetic progression $n\equiv 0\,(\operatorname{mod}{d})$: There are an appropriate approximation $\mathfrak{M}$ to $A_{1}$ and a multiplicative function $g$ supported on squarefree numbers with all its prime factors belonging to $\mathcal{P}$ verifying
$0<g(p)<1\quad(p\in\mathcal{P})$
such that
(a) the remainder
$r(\mathcal{A},d)=A_{d}-g(d)\mathfrak{M}$
is small on average over $d\mid P(z)$;
(b) there exists a constant $L>1$ such that
for $2\leqslant z_{1}<z_{2}$.
Let $F$ and $f$ be the continuous solutions to the system
where $\gamma$ is the Euler constant.
Let $r\geqslant 2$ be a positive integer. We say that $\lambda$ is well-factorable of degree $r$, if for every decomposition $D=D_{1}D_{2}\cdots D_{r}$ with $D_{1},D_{2},\dots,D_{r}\geqslant 1,$ there exist $r$ arithmetic functions $\lambda_{1},\lambda_{2},\dots,\lambda_{r}$ such that
$\lambda=\lambda_{1}*\lambda_{2}*\cdots*\lambda_{r}$
with each $\lambda_{j}$ of level $D_{j}$ and order $k$.
We now state the following fundamental result of Iwaniec [Iw80].
Lemma 2.1.
Let $r\geqslant 2$ and $0<\varepsilon<\tfrac{1}{8}$. Under the above hypothesis, we have
$S(\mathcal{A},\mathcal{P},z)\leqslant\mathfrak{M}V(z)\bigg{\{}F\bigg{(}\frac{\log D}{\log z}\bigg{)}+E\bigg{\}}+\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}(d)r(\mathcal{A},d)$
for all $z\geqslant 2$, where $F$ is given by the system (2.1), $T$ depends only on $\varepsilon$, $\lambda_{t}(d)$ is well-factorable of level $D,$ degree $r$ and order $1$, and $E\ll\varepsilon+\varepsilon^{-8}\mathrm{e}^{L}(\log D)^{-1/3}.$
As one can show, $F(s)$ decays rapidly to $1$ as $s\rightarrow+\infty.$ In applications, we would like to explore the admissible $D$ as large as possible such that the total remainder
$\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}(d)r_{d}(\mathcal{A})$
is still under control. In principle, a larger $D$ would produce a better upper bound for $S(\mathcal{A},\mathcal{P},z)$ upon a suitable choice for $z.$ The merit of Lemma 2.1 lies in the well-factorization of $\lambda_{t}$, so that one can estimate the resultant bilinear forms more effectively to produce a better admissible $D$ (compared with earlier sieve methods before Iwaniec). This bilinear structure is very crucial in many existing applications of Iwaniec’s linear sieve, and in fact only bilinear structure is employed in such applications. As one may see from Lemma 2.1, we can take $r$ to be a large integer, so that some multilinear structure appears, and a natural question is whether we can utilize this new structure effectively in practice. As an instance of this feature, we need to express $\lambda_{t}$, in the proof of Theorem 1.1 for prime moduli, as a convolution of four arithmetic functions.
@@ 3. Dirichlet polynomials and character sums
Let $F\in\mathcal{C}_{0}^{\infty}(\mathbf{R}^{2})$ be a Schwartz function with support in $]N,2N]\times]N,2N]$, satisfying
$\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial x_{1}^{\nu_{1}}\partial x_{2}^{\nu_{2}}}F(x_{1},x_{2})$ $\ll\eta^{\nu_{1}+\nu_{2}}N^{-\nu_{1}-\nu_{2}}$
@@
We have the following asymptotic formula.
@@
(3.2) $\mathcal{R}$ $=\frac{\varphi(q)}{q^{2}}\Big{(}\mathop{\sum\sum}_{\begin{subarray}{c}m_{1},m_{2}\sim M\\ (m_{1}m_{2},q)=1\end{subarray}}\alpha_{m_{1}}\beta_{m_{2}}\Big{)}\Big{(}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z\Big{)}+\mathcal{E},$
@@
$\mathcal{E}\ll\eta^{\frac{9}{2}}M^{\frac{3}{2}}Nq^{\frac{\vartheta-1}{2}}(1+Mq^{-\frac{1}{2}})^{\frac{3}{2}}(MNq)^{\varepsilon}+q^{-1}(MN)^{2}(\log MNq)^{-A}+\eta(MN)^{1+\varepsilon}$
for any $\varepsilon,A>0,$ with an implied constant depending only on $(\varepsilon,A)$. Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
@@
Let $C,M,N\geqslant 1,$ $a,r,s\in\mathbf{Z}^{+}$ pairwise coprime, and $g(c,m,n)$ a smooth function with compact support in $[C,2C]\times[M,2M]\times[N,2N]$ such that
@@
Theorem 3.1 allows us to derive the following version of Lindelöf Hypothesis on average.
@@
Let $q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=(\lambda_{n})_{n\leqslant N}$ be a well-factorable function of level $N.$ Then we have
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n}\lambda_{n}\chi(n)\Big{|}^{2}\Big{|}\sum_{l\in\mathbf{Z}}g\Big{(}\frac{l}{L}\Big{)}\chi(l)\Big{|}^{2}$
$\ll\eta^{\frac{9}{2}}N^{\frac{3}{2}}Lq^{\frac{\vartheta-1}{2}}(1+Nq^{-\frac{1}{2}})^{\frac{3}{2}}q^{\varepsilon}+\Big{(}\eta^{\frac{9}{2}}+\frac{NL}{q}\Big{)}NL(\log q)^{-A}$
@@
The following lemma is an unconditional case of Conjecture 1.1, provided that the Dirichlet polynomial is not too long.
@@
Let $q$ be a large prime and $N\leqslant q^{\frac{1}{4}}$. Let $\boldsymbol{\alpha}=(\alpha_{n})_{n\leqslant N}$ be an arbitrary coefficient supported on squarefree numbers. Then we have
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n\leqslant N}\alpha_{n}\chi(n)\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll(1+|t|)^{12}Nq^{\varepsilon}\|\boldsymbol{\alpha}\|_{\infty}$
@@
As stated in Proposition 1.1, Lemma 3.4 is proved by Blomer, Humphries, Khan and Milinovich [BHKM20] for $t=0$, and their argument can be easily extended to all $t\in\mathbf{R}$. It is very useful to relax $q$ to an arbitrary modulus, which in turn can guarantee Theorem 1.1 with $\vartheta=0$ for all $q\geqslant 3$.
@@ 4.2. Composite trace functions and arithmetic exponent pairs
By the Chinese remainder theorem, a trace function can be also defined for squarefree moduli, which allows Wu and Xi [WX21] to develop the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs.
@@
$A^{k}B\cdot(0,1,0)$ $=\Big{(}\frac{1}{2^{k+1}-2},1-\frac{k}{2^{k+1}-2},\frac{1}{2^{k+1}-2}\Big{)}$
for each positive integer $k\geqslant 2$.
@@
$c=\frac{1}{2^{k+1}-2}\Big{(}k-1-\frac{1}{\theta}\Big{)}.$
@@
5. Proofs of Theorems 1.1, 1.5
5.1. Initial treatment by linear sieves
@@
In view of the definition of $C(\varpi)$ in (1.3), the above error $O(q^{-1}x(\log x)^{-2})$ is acceptable, and it suffices to derive upper bounds for the smooth sum $\pi^{*}(x;q,a).$
@@
$\pi^{*}(x;q,a)\leqslant\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}A(d)\leqslant\frac{x}{q}V(z)F\Big{(}\frac{\log D}{\log z}\Big{)}+\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}r(d),$
@@
Thanks to Lemma 2.1, we may choose well-factorable remainder terms in the above application of linear Rosser–Iwaniec sieve. We thus conclude the following result.
@@
Let $\varepsilon>0$ and $D<x^{1-\varepsilon}.$ We have
@@
(5.4) $R(D):=\sum_{d\leqslant D}\lambda(d)r(d)\ll\frac{x}{q(\log x)^{2024}}$
with $D$ as large as possible.

### v3

@@ Acknowledgements
We are very grateful to Étienne Fouvry and Kaisa Matomäki for their kind comments and suggestions. This work is supported in part by NSFC (No. 12025106).
@@ 2. Dirichlet polynomials and character sums
Let $F\in\mathcal{C}_{0}^{\infty}(\mathbf{R}^{2})$ be a Schwartz function with support in $]L,2L]\times]L,2L]$, satisfying
$\frac{\partial^{\nu_{1}+\nu_{2}}}{\partial x_{1}^{\nu_{1}}\partial x_{2}^{\nu_{2}}}F(x_{1},x_{2})$ $\ll\eta^{\nu_{1}+\nu_{2}}L^{-\nu_{1}-\nu_{2}}$
@@
We have an asymptotic formula for the above counting function.
Theorem 2.1.
Let $\eta,q,N,L\geqslant 1.$ Suppose $-β\boldsymbol{\alpha},\bbeta$ are complex coefficients supported in $[N,2N],$ and $F$ is a Schwartz function defined as above. Then
$\mathcal{R}$ $=\frac{\varphi(q)}{q^{2}}\Big{(}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\alpha_{n_{1}}\beta_{n_{2}}\Big{)}\Big{(}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z\Big{)}$
$\ \ \ \ \ +O(\eta(NL)^{1+\varepsilon}+\eta^{\frac{7}{2}}N^{\frac{23}{8}}Lq^{-1+\varepsilon}(N^{-\frac{9}{40}}q^{\frac{3}{20}}+1))$
for any $\varepsilon>0,$ with an implied constant depending only on $\varepsilon.$
Theorems 2.1 and 2.2 will be utilized to study a weighted moment of character sums, which can be regarded as a kind of Lindelöf Hypothesis on average. The proof of Theorem 2.1 relies on the following estimate, due to Bettin and Chandee [BC18], for trilinear forms involving Kloosterman fractions.
Lemma 2.3.
[…]
×$\qquad\times((M+N)^{\frac{1}{4}}+(KMN)^{\frac{1}{40}}(KM+KN)^{\frac{1}{8}}).$
Remark 1.
The above estimate of Bettin and Chandee [BC18] is in fact a successor of a (weaker) bound due to Duke, Friedlander and Iwaniec [DFI97b] with $K=1$.
If the coefficients $-β\boldsymbol{\alpha},\bbeta$ have suitable convolution structures, we can do a bit better than Theorem 2.1.
@@
(2.2) $\mathcal{R}$ $=\frac{\varphi(q)}{q^{2}}\Big{(}\mathop{\sum\sum}_{\begin{subarray}{c}n_{1},n_{2}\sim N\\ (n_{1}n_{2},q)=1\end{subarray}}\alpha_{n_{1}}\beta_{n_{2}}\Big{)}\Big{(}\iint_{\mathbf{R}^{2}}F(y,z)\mathrm{d}y\mathrm{d}z\Big{)}+\mathcal{E},$
@@
$\mathcal{E}$ $\ll\eta^{\frac{9}{2}}N^{\frac{3}{2}}Lq^{\frac{\vartheta-1}{2}}(N_{1}N_{2})^{\frac{1}{2}}(N^{2}q^{-1}+N_{1}N_{2})^{\frac{1}{4}}\Big{(}1+\frac{N^{2}}{qN_{1}N_{2}}\Big{)}^{\frac{1}{4}}(qNL)^{\varepsilon}$
$\ \ \ \ +\eta(NL)^{1+\varepsilon}+\eta^{\frac{9}{2}}N^{2}L(qN_{1}N_{2})^{-\frac{1}{2}}(qNL)^{\varepsilon}$
for any $\varepsilon,A>0,$ with an implied constant depending only on $(\varepsilon,A,c)$. Here $\vartheta\leqslant 7/64$ is the exponent towards the Ramanujan–Petersson conjecture for $\mathrm{GL}_{2}(\mathbf{Q}).$
Remark 2.
At first glance, the restriction that $\boldsymbol{\alpha}^{\prime\prime}$ and $-β′′\bbeta^{\prime\prime}$ vanish on the set $\{n\geqslant 2:\omega(n)>(\log N)^{1/3}\}$ is quite artificial. In our applications (as well as many other situations) this is harmless since only a minority of integers have been ruled out. We remark that this condition is designed for the later application of a trick due to Fouvry (see Lemma 7.1 below).
@@
Let $C,M,N\geqslant 1,$ $a,r,s\in\mathbf{Z}^{+}$ pairwise coprime with $\mu(rs)\neq 0,$ and $g(c,m,n)$ a smooth function with compact support in $[C,2C]\times[M,2M]\times[N,2N]$ such that
@@
Theorems 2.1 and 2.2 allow us to derive the following versions of Lindelöf Hypothesis on average.
@@
Let $N,L,q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=(\lambda_{n})$ be an arbitrary coefficient with support in $[1,N].$ Suppose that $N,L\ll q^{O(1)}.$ Then we have
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n}\frac{\lambda_{n}\chi(n)}{\sqrt{n}}\Big{|}^{2}\Big{|}\sum_{l\in\mathbf{Z}}g\Big{(}\frac{l}{L}\Big{)}\frac{\chi(l)}{\sqrt{l}}\Big{|}^{2}\ll\eta q^{\varepsilon}+\eta^{\frac{7}{2}}N^{\frac{15}{8}}q^{-1+\varepsilon}(N^{-\frac{9}{40}}q^{\frac{3}{20}}+1)$
for any $\varepsilon>0,$ where the implied constant depends only on $\varepsilon.$
Lemma 2.6.
Let $N_{1},N_{2},N,L,q\geqslant 3$ and $g$ a smooth function with compact support in $[1,2],$ satisfying $\|g^{(j)}\|_{\infty}\ll\eta^{j}$ for some $\eta\geqslant 1.$ Let $\boldsymbol{\lambda}=\boldsymbol{\lambda}_{1}*\boldsymbol{\lambda}_{2}$ be a convolution, where $\boldsymbol{\lambda}_{1},$ $\boldsymbol{\lambda}_{2}$ have supports in $[1,N_{1}],$ $[1,N_{2}]$ with $N_{1}N_{2}=N$ and $\|\boldsymbol{\lambda}_{1}\|_{\infty},\|\boldsymbol{\lambda}_{2}\|_{\infty}\leqslant 1.$ We also assume that $\boldsymbol{\lambda}$ vanishes on the set $\{n\geqslant 2:\omega(n)>(\log N)^{1/3}\}.$ Suppose that $N,L\ll q^{O(1)}.$ Then we have
$\frac{1}{\varphi(q)}$ $\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n}\frac{\lambda_{n}\chi(n)}{\sqrt{n}}\Big{|}^{2}\Big{|}\sum_{l\in\mathbf{Z}}g\Big{(}\frac{l}{L}\Big{)}\frac{\chi(l)}{\sqrt{l}}\Big{|}^{2}$
$\ll\eta^{\frac{9}{2}}q^{\frac{\vartheta-1}{2}}N_{1}^{\frac{1}{2}}N_{2}^{2}(N_{1}q^{-\frac{1}{2}}+1)q^{\varepsilon}+\eta q^{\varepsilon}+\eta^{\frac{9}{2}}N_{1}q^{-\frac{1}{2}+\varepsilon}$
@@
The following lemma presents a weighted fourth moment of Dirichlet $L$-functions.
@@
Let $q$ be a large prime and $N\geqslant 1$. Let $\boldsymbol{\alpha}=(\alpha_{n})_{n\leqslant N}$ be an arbitrary coefficient supported on squarefree numbers and $\|\boldsymbol{\alpha}\|_{\infty}\leqslant 1.$ Then we have
$\frac{1}{\varphi(q)}\sum_{\begin{subarray}{c}\chi\,(\operatorname{mod}{q})\\ \chi\neq\chi_{0}\end{subarray}}\Big{|}\sum_{n\leqslant N}\frac{\alpha_{n}\chi(n)}{\sqrt{n}}\Big{|}^{2}|L(\tfrac{1}{2}+it,\chi)|^{4}\ll(1+|t|)^{16}(N^{2}q^{-\frac{1}{2}}+1)q^{\varepsilon}$
@@
Lemma 2.7 is proved by Blomer, Humphries, Khan and Milinovich [BHKM20] for $t=0$, and their argument can be easily extended to all $t\in\mathbf{R}$. It is very useful to relax $q$ to an arbitrary modulus, which in turn can guarantee the choice (1.6) in Theorem 1.1 for all $q\geqslant 3$.
@@ 3.2. Composite trace functions and arithmetic exponent pairs
By the Chinese remainder theorem, a trace function can be also defined for squarefree moduli, which allows Wu, Xi and Sawin [WX21] to develop the general form of $q$-analogue of van der Corput method and arithmetic exponent pairs.
@@
$A^{k-1}B\cdot(0,1,0)$ $=\Big{(}\frac{1}{2^{k+1}-2},1-\frac{k}{2^{k+1}-2},\frac{1}{2^{k+1}-2}\Big{)}$
for each positive integer $k\geqslant 3$.
@@
$c=\frac{1}{2^{k+1}-2}\Big{(}k+1-\frac{1}{\theta}\Big{)}.$
@@
Remark 3.
We would like to mention that the statement in Lemma 3.6 is not new. One may refer to Graham and Ringrose [GR90, Theorem 5] for upper bounds with explicit dependences on the factorization of $q$.
@@
4. Sieve methods
4.1. Linear sieve
$A_{d}=\sum_{n\equiv 0\!\,(\operatorname{mod}{d})}a_{n}.$
In applications to sieve methods, it is expected that $\mathcal{A}$ equidistributes in the special arithmetic progression $n\equiv 0\,(\operatorname{mod}{d})$: There are an appropriate approximation $\mathscr{X}$ to $A_{1}$ and a multiplicative function $g$ supported on squarefree numbers with all its prime factors belonging to $\mathcal{P}$ verifying
$0<g(p)<1\quad(p\in\mathcal{P})$
such that
(a) the remainder
$r(\mathcal{A},d)=A_{d}-g(d)\mathscr{X}$
is small on average over $d\mid P(z)$;
(b) there exists a constant $L>1$ such that
for $2\leqslant z_{1}<z_{2}$.
Let $F$ and $f$ be the continuous solutions to the system
where $\gamma\approx 0.57721$ is the Euler constant.
We say that $\lambda$ is well-factorable, if for every decomposition $D=D_{1}D_{2}$ with $D_{1},D_{2}\geqslant 1,$ there exist two arithmetic functions $\lambda_{1},\lambda_{2}$ such that
$\lambda=\lambda_{1}*\lambda_{2}$
with each $\lambda_{j}$ of level $D_{j}$ and order $k$.
We now state the following fundamental result of Iwaniec [Iw80].
Lemma 4.1.
Let $0<\varepsilon<\tfrac{1}{8}$. Under the above hypothesis, we have
$S(\mathcal{A},\mathcal{P},z)\leqslant\mathscr{X}V(z)\bigg{\{}F\bigg{(}\frac{\log D}{\log z}\bigg{)}+E^{+}\bigg{\}}+\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{+}(d)r(\mathcal{A},d)$
and
$S(\mathcal{A},\mathcal{P},z)\geqslant\mathscr{X}V(z)\bigg{\{}f\bigg{(}\frac{\log D}{\log z}\bigg{)}-E^{-}\bigg{\}}-\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{-}(d)r(\mathcal{A},d)$
for all $z\geqslant 2,$ where $F,f$ are given by the system (4.1), $T$ depends only on $\varepsilon,$ $\lambda_{t}^{\pm}(d)$ are well-factorable of level $D$ and order $1,$ and $E^{\pm}\ll\varepsilon+\varepsilon^{-8}\mathrm{e}^{L}(\log D)^{-1/3}.$
From the system (4.1), one can show that
(4.2) $F(s)=\begin{cases}\displaystyle\frac{2\mathrm{e}^{\gamma}}{s},&\text{for }1\leqslant s\leqslant 3,\\ \vskip 2.84526pt\cr\displaystyle\frac{2\mathrm{e}^{\gamma}}{s}\Big{(}1+\int_{2}^{s-1}\frac{\log(t-1)}{t}\mathrm{d}t\Big{)},&\text{for }3\leqslant s\leqslant 5,\end{cases}$
and
(4.3) $f(s)=\begin{cases}0,&\text{for }0\leqslant s\leqslant 2,\\ \vskip 2.84526pt\cr\displaystyle\frac{2\mathrm{e}^{\gamma}}{s}\log(s-1),&\text{for }2\leqslant s\leqslant 4.\end{cases}$
Moreover, $F(s)$ decreases (resp. $f(s)$ increases) rapidly to $1$ as $s\rightarrow+\infty.$ In applications, we would like to explore the admissible $D$ as large as possible such that the total remainder
$\sum_{t\leqslant T}\sum_{d\mid P(z)}\lambda_{t}^{\pm}(d)r_{d}(\mathcal{A})$
is still under control. In principle, a larger $D$ would produce a better upper/lower bound for $S(\mathcal{A},\mathcal{P},z)$ upon a suitable choice for $z.$ The merit of Lemma 4.1 lies in the well-factorization of $\lambda_{t}^{\pm}$, so that one can estimate the resultant bilinear forms more effectively to produce a better admissible $D$ (compared with earlier sieve methods before Iwaniec). This bilinear structure is very crucial in many existing applications of Iwaniec’s linear sieve.
We also quote the following Buchstab identity, which can be verified directly.
Lemma 4.2.
For all $z>w\geqslant 2,$ we have
$S(\mathcal{A},\mathcal{P},z)=S(\mathcal{A},\mathcal{P},w)-\sum_{\begin{subarray}{c}w\leqslant p<z\\ p\in\mathcal{P}\end{subarray}}S(\mathcal{A}_{p},\mathcal{P},p).$
4.2. Sifting primes in $\pi(x,q;a)$
@@
In view of the definition of $C(\varpi)$ in (1.3), the above error $O(q^{-1}x(\log x)^{-2})$ is acceptable, and it suffices to derive upper bounds for the smooth sum $\pi^{*}(x;q,a).$ Denote by $P(t)$ the product of primes up to $t$ and not dividing $q$. On one hand, we have
(4.4) $\pi^{*}(x;q,a)\leqslant\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ (n,P(z))=1\end{subarray}}\varPhi(n)+O(1+q^{-1}x^{\frac{1}{2}})$
for all $2\leqslant z<\sqrt{x}.$ By Buchstab identity (Lemma 4.2), we also have
(4.5) $\pi^{*}(x;q,a)=\sum_{\begin{subarray}{c}n\equiv a\,(\operatorname{mod}{q})\\ (n,P(w))=1\end{subarray}}\varPhi(n)-\sum_{w\leqslant p<z}\sum_{\begin{subarray}{c}pn\equiv a\,(\operatorname{mod}{q})\\ (n,P(p))=1\end{subarray}}\varPhi(pn)+O(1+q^{-1}x^{\frac{1}{2}})$
for all $2\leqslant w<z<\sqrt{x}.$
@@
(4.8) $\pi^{*}(x;q,a)$ $\leqslant\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}A(d)+O(q^{-1}x(\log x)^{-2})$
(4.9) $\leqslant\frac{x}{q}V(z)F\Big{(}\frac{\log D}{\log z}\Big{)}+\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(z)\end{subarray}}\lambda_{d}r(d)+O(q^{-1}x(\log x)^{-2}),$
@@
Thanks to Lemma 4.1, we may choose well-factorable remainder terms in the above application of the linear Rosser–Iwaniec sieve. We thus conclude the following result.
@@
Let $\varepsilon>0$ sufficiently small and $D<x^{1-\varepsilon}.$ For any $z$ with $D^{\frac{1}{3}}\leqslant z<\sqrt{x},$ we have
@@
In place of (4.4), we may obtain the following upper bound for $\pi(x;q,a)$ by virtue of (4.5).
Proposition 4.2.
Let $\varepsilon>0$ sufficiently small and $D<x^{1-\varepsilon}.$ For any $2\leqslant w<z<\sqrt{x},$ we have
$\pi(x;q,a)$ $\leqslant(\mathrm{e}^{-\gamma}+O(\varepsilon))\frac{x}{\varphi(q)}\Big{\{}\frac{1}{\log w}F\Big{(}\frac{\log D}{\log w}\Big{)}-\sum_{w\leqslant p<z}\frac{1}{p\log p}f\Big{(}\frac{\log D_{p}}{\log p}\Big{)}\Big{\}}$
$\ \ \ \ +\sum_{t\leqslant T^{+}(\varepsilon)}\sum_{\begin{subarray}{c}d\leqslant D\\ d\mid P(w)\end{subarray}}\lambda_{t}^{+}(d)r(d)+\sum_{t\leqslant T^{-}(\varepsilon)}\sum_{w\leqslant p<z}\sum_{\begin{subarray}{c}d\leqslant D_{p}\\ d\mid P(p)\end{subarray}}\lambda_{t}^{-}(d)r(pd),$
where $T^{\pm}(\varepsilon)$ depends only on $\varepsilon,$ and $\lambda_{t}^{+},\lambda_{t}^{-}$ are well-factorable of level $D,D_{p},$ respectively.
5. Proofs of Theorems 1.1 and 1.5
@@
(5.1) $R^{+}(D):=\sum_{d\leqslant D}\lambda^{+}(d)r(d)\ll\frac{x}{q(\log x)^{2024}}$
and
(5.2) $R^{-}(P,D^{*}):=\sum_{p\sim P}\sum_{d\leqslant D^{*}}\lambda^{-}(d)r(pd)\ll\frac{x}{q(\log x)^{2024}}$
with $D,D^{*}$ as large as possible.